OpenStudy (anonymous):

In a certain county, the number of charter schools is 6 less than twice the number of alternative schools. We know that there are 52 charter schools in the county. How many alternative schools are in the county?

OpenStudy (anonymous):

@nevermind_justschool @nincompoop

OpenStudy (adamaero):

What equation are you thinking, to begin?

OpenStudy (anonymous):

C = 2A - 6?

OpenStudy (anonymous):

@adamaero

OpenStudy (adamaero):

What can you do with the 52 charter schools?

OpenStudy (adamaero):

Do you need another equation?

OpenStudy (anonymous):

I'm just not sure how i would do that

OpenStudy (anonymous):

like after that

OpenStudy (adamaero):

Where could you put the 52? Guess if you are not 100% sure.

OpenStudy (aaronandyson):

x = 2y - 6 52 = 2y-6

OpenStudy (aaronandyson):

Taking 52 = 2y - 6 We add 6 on both sides 52+6 = 26 + 6 -6 The two 6 on the right hand side cancel out to become 0 58 = 2y Divide both sides by 2 58/2 = 2y/2 29 = y

OpenStudy (aaronandyson):

y = 29 x = 2y-6 x = 2(29) - 6 x = 58 - 6 x = 52